### v1

@@ 1 Introduction
In recent years, Continual Learning (CL) [15, 16, 17] emerges as a conceptual solution for addressing CF in AI systems. CL aims to balance the system’s ability to learn new tasks without forgetting the previous knowledge, known as the stability-plasticity dilemma [18, 19] (Fig. 1). It also seeks to achieve generalizability across tasks. In addition to avoiding CF, the primary objectives of CL also include ensuring scalability of NNs, minimizing reliance on old data, realizing controlled forgetting, and enabling rapid adaptation and recovery, which will be further discussed in Section 2.2.
@@
1.1 Surveys on Continual Learning (CL)
@@
The remarkable CL capability of humans has sparked interest in the emerging of NCL paradigm, i.e., investigation on how SNN architectures and learning rules can be utilized to develop efficient CL systems. However, notable surveys on NCL still focused on exploring biological mechanisms for CL [34], and providing the background theories [54]. Therefore, they have not provided a holistic survey that provides a comprehensive view of NCL, ranging from theories to applications. Toward this, in this paper, we provide an in-depth review on preliminaries of CL and NC, state-of-the-art advancements of CL research in ANNs/DNNs and SNNs, and real-world applications for NCL (Fig. 2). Moreover, we also discuss the open research challenges, hence offering an updated perspective for future research directions in developing energy-efficient and adaptive AI systems. A comparison summary of our survey and the existing ones is presented in Table 1.
@@
Table 1: A qualitative comparison of our survey and the existing ones. Note: “✓”, “+”, and “-” mean full, partial, and no consideration, respectively.
@@ 1.2 Our Novel Contributions
1.3 Paper Structure
The rest of this survey paper is organized as shown in Fig. 3. Section 2 delves into the CL fundamentals. Section 3 provides an overview of SNNs and the state-of-the-art NCL methods. In Section 4, a summary of real-world application use-cases of NCL is presented. Then, Section 5 discusses open research challenges for meeting the CL desiderata with SNNs. Finally, Section 6 presents the conclusion.
2 Continual Learning
@@ 2.2.1 Scalability
2.2.2 No/Minimal Usage of Old Data
@@ 2.2.3 Task Agnostic (TA)
2.2.4 Positive Forward Transfer (FWT)
@@
2.2.5 Positive Backward Transfer (BWT)
@@ 2.2.6 Controlled Forgetting
2.2.7 Fast Adaptation and Recovery
@@
2.3 Scenarios in CL
CL scenarios can be categorized using a taxonomy presented in Fig. 5, while their formal comparison is provided in Table 2. The initially-identified scenarios are Task-Incremental Learning (Task-IL), Domain-Incremental Learning (Domain-IL), and Class-Incremental Learning (Class-IL). While these scenarios cover some desiderata of CL, they fail to meet the requirement of TA [53]; see Fig. 6. They assume that tasks have clear and well-defined boundaries during training. Therefore, recent studies proposed expanding the CL scenarios by introducing Discrete TA (DTA) [74], Continuous TA (CTA) [74], and Open-World TA (OWTA) [53].
@@ 2.3 Scenarios in CL
CL Scenario Difference between $\mathcal{D}_{t-1}$ and $\mathcal{D}_{t}$ Task ID Online Learning
p$(\mathcal{X}_{t-1})\neq$
p$(\mathcal{X}_{t})$
p$(\mathcal{Y}_{t-1})\neq$
p$(\mathcal{Y}_{t})$
@@
2.3.1 Task-Incremental Learning (Task-IL)
@@
2.3.2 Domain-Incremental Learning (Domain-IL)
2.3.3 Class-Incremental Learning (Class-IL)
2.3.4 Discrete Task Agnostic (DTA)
@@
2.3.5 Open-World Task Agnostic (OWTA)
2.3.6 Continuous Task Agnostic (CTA)
[…]
2.4 Online Continual Learning (OCL)
@@ 2.5 CL Methods
2.5.1 Representation-based Approach
@@
Figure 7: A visual illustration of the taxonomy of (a) representation-based approach and (b) its respective methods; adapted from studies in [16].
  1. A.
@@
2.5.2 Regularization-based Approach
@@
Figure 8: A visual illustration of the taxonomy of (a) regularization-based approach and (b) its respective methods; adapted from studies in [16].
  1. A.
@@
2.5.3 Rehearsal/Replay-based Approach
@@
Figure 9: A visual illustration of the taxonomy of (a) replay-based approach and (b) its respective methods; adapted from studies in [16].
  1. A.
@@
2.5.4 Optimization-based Approach
@@
Figure 10: A visual illustration of the taxonomy of (a) optimization-based approach and (b) its respective methods; adapted from studies in [16].
  1. A.
@@
2.5.5 Architecture-based Approach
@@
Figure 11: A visual illustration of the taxonomy of (a) architecture-based approach and (b) its respective methods; adapted from studies in [16].
  1. A.
@@
    Figure 12: A visual illustration of the model decomposition method exhibiting two types; corresponding to parameters (i.e., low-rank factorization) and representations (i.e., masks of intermediate features); adapted from studies in [16].
All these existing methods for fulfilling the desiderata of CL are also summarized and presented in Table 3.
Table 3: Summary of CL methods based on Representation-, Regularization-, Rehearsal/Replay- and Architecture-based approaches that potentially fulfill “+” or have fulfilled the desiderata of CL.
Approach Methods Work,(Year) Desiderata of CL
@@
[…]
2.6 CL Settings
2.6.1 Supervised Learning
@@ 2.7 Evaluation Metrics
3 Neuromorphic Computing
3.1 Overview of SNNs
@@
3.2 Neuromorphic Continual Learning (NCL)
NCL is an advancement of bio-plausible SNNs for facilitating continual adaptation in dynamic environments. Here, we discuss the state-of-the-art NCL methods, including their key ideas and limitations, whose details are summarized in Table 4.
3.2.1 Enhancements on Unsupervised STDP Learning
While conventional STDP enables efficient online unsupervised learning, employing STDP alone may still suffer from CF [41, 42]. To address these challenges, proposed techniques have been proposed in the literature, as discussed below.
@@ 3.2.3 Active Dendrites
3.2.4 Bayesian Continual Learning
@@
3.2.5 Architecture-based Approach
@@
3.2.6 Rehearsal/Replay-based Approach
@@
3.2.7 Regularization-based Approach
@@
Figure 24: Illustration of the HLOP; adapted from studies in [319]. (a) Overview of the method. (b) Skew-symmetric connection weights in lateral circuits. (c) Construction of new subspaces for new tasks with Hebbian learning in lateral circuits. (d) Recurrent lateral connections based orthogonal projection.
@@ 3.2.8 Hebbian Learning
3.3 Efficiency Enhancement Methods for NCL
@@ 3.3.2 Weight Quantization
Table 4: Summary of the state-of-the-art works for energy-efficient CL employing bio-plausible SNNs (i.e., NCL).
@@
Table 5: Comparative analysis of the state-of-the-art in terms of network architecture, model size, performance, memory footprint, and power/energy savings. Note, FC: fully-connected layer, and CONV: convolutional layer.
@@
3.4 Summary of NCL
The state-of-the-art methods have performed initial studies for enabling NCL (Table 4). We observe that methods with bio-plausible learning rules (e.g., STDP) facilitate unsupervised learning, which is suitable for OCL scenario. These works mainly employ fully-connected SNN architectures (see Table 5), consisting of input layer and a pair of excitatory-inhibitory layer. As the input data grows in complexity, fully-connected SNNs struggle to capture the important hierarchical features without increasing the network size. Larger architectures often lead to better accuracy but require larger storage sizes, which may exceed the resource constraints. This limits the deployable networks on practical hardware platforms, and make their performances limited to simple datasets (e.g., MNIST) with simple CL scenarios (e.g., Split-MNIST). Meanwhile, methods that employ supervised-based learning rules (e.g., surrogate gradient-based backpropagation) can achieve acceptable performance on complex datasets (e.g., ImageNet). However, they are not suitable for OCL scenario, as they require costly labeled data and power-hungry training. Furthermore, these works mainly focus on energy-efficiency, and have not optimized resource utilization. Therefore, alternative methods are required especially for enabling energy-efficient SNNs with OCL capabilities, which are beneficial for many real-world application use-cases.
4 Real-World Application Use-Cases
@@
4.1 Adaptive Robots
4.1.1 Object Recognition for Humanoid Robots
@@
4.1.2 Inverse Kinematics and PID Control
@@
4.2 Autonomous Vehicles and Mobile Agents
@@ 4.2.1 Cars Detection
4.2.2 Road Lane Detection
Other SNN works for low-power autonomous vehicles include the road line detection. For instance, LaneSNN [348] addressed the imbalance between lane and background classes by employing a novel loss function combining Weighted Binary Cross Entropy (WCE) and Mean Squared Error (MSE); see Table 6. Evaluations demonstrated comparable performance to state-of-the-art methods like LaneNet and RefineNet but with significantly lower power consumption. However, it still considered supervised settings which require costly labeled data, power-hungry training, and long training time to update its knowledge, hence it is not suitable for OCL scenario.
@@ 4.2.2 Road Lane Detection
4.2.3 Simultaneous Localization and Mapping (SLAM)
@@ 4.2.4 People Detection
A summary of real-world application use-cases that will benefit from SNNs with OCL, is presented in Table 6.
Table 6: Summary of real-world application use-cases that will benefit from SNNs with OCL capabilities, and their current state-of-the-art works.
Work CL Dataset CL Learning Neuron Implementation
Applications (Year) Setting Approach Rule Model Hardware
Adaptive Robots Object [341] Supervised Custom Representation- Modified LIF Loihi
@@
SLAM [349] Unsupervised Real-time data Representation- STDP LIF n/a
@@
[…]
5 Open Research Challenges
@@
Adapting the Existing CL Methods in SNN Domain: Investigating how existing CL methods in Section 2.5 can be adapted for SNNs is potential for combining their strengths. For instance, adapting the regularization-based methods and how it impacts the performance of SNNs trained with local learning rules, particularly in dynamic environments could be explored. The optimal memory size for replay-based method in SNNs that balances desiderata effectively, particularly in resource-constrained environments can also be explored.
@@ 6 Conclusion
Acknowledgements
This work was supported by United Arab Emirates University (UPAR) with Fund code 12N169. This work was also partially supported by the NYUAD Center for Interacting Urban Networks (CITIES), funded by Tamkeen under the NYUAD Research Institute Award CG001, and the NYUAD Center for Artificial Intelligence and Robotics (CAIR), funded by Tamkeen under the NYUAD Research Institute Award CG010.

### v2

@@ I Introduction
In recent years, Continual Learning (CL) [15, 16, 17] emerges as a conceptual solution for addressing CF in AI systems. CL aims to balance the system’s ability to learn new tasks without forgetting the previous knowledge, known as the stability-plasticity dilemma [18, 19] (Fig. 1). It also seeks to achieve generalizability across tasks. In addition to avoiding CF, the primary objectives of CL also include ensuring scalability of NNs, minimizing reliance on old data, realizing controlled forgetting, and enabling rapid adaptation and recovery, which will be further discussed in Section II-B.
@@
I-A Surveys on Continual Learning (CL)
@@
The remarkable CL capability of humans has sparked interest in the emerging of NCL paradigm, i.e., investigation on how SNN architectures and learning rules can be utilized to develop efficient CL systems. However, notable surveys on NCL still focused on exploring biological mechanisms for CL [34], and providing the background theories [54]. Therefore, they have not provided a holistic survey that provides a comprehensive view of NCL, ranging from theories to applications. Toward this, in this paper, we provide an in-depth review on preliminaries of CL and NC, state-of-the-art advancements of CL research in ANNs/DNNs and SNNs, and real-world applications for NCL (Fig. 2). Moreover, we also discuss the open research challenges, hence offering an updated perspective for future research directions in developing energy-efficient and adaptive AI systems. A comparison summary of our survey and the existing ones is presented in Table I.
@@
TABLE I: A qualitative comparison of our survey and the existing ones. Note: “✓”, “+”, and “-” mean full, partial, and no consideration, respectively.
@@ I-B Our Novel Contributions
I-C Paper Structure
The rest of this survey paper is organized as shown in Fig. 3. Section II delves into the CL fundamentals. Section III provides an overview of SNNs and the state-of-the-art NCL methods. In Section IV, a summary of real-world application use-cases of NCL is presented. Then, Section V discusses open research challenges for meeting the CL desiderata with SNNs. Finally, Section VI presents the conclusion.
II Continual Learning
@@ II-B1 Scalability
II-B2 No/Minimal Usage of Old Data
@@ II-B3 Task Agnostic (TA)
II-B4 Positive Forward Transfer (FWT)
@@
II-B5 Positive Backward Transfer (BWT)
@@ II-B6 Controlled Forgetting
II-B7 Fast Adaptation and Recovery
@@
II-C Scenarios in CL
CL scenarios can be categorized using a taxonomy presented in Fig. 5, while their formal comparison is provided in Table II. The initially-identified scenarios are Task-Incremental Learning (Task-IL), Domain-Incremental Learning (Domain-IL), and Class-Incremental Learning (Class-IL). While these scenarios cover some desiderata of CL, they fail to meet the requirement of TA [53]; see Fig. 6. They assume that tasks have clear and well-defined boundaries during training. Therefore, recent studies proposed expanding the CL scenarios by introducing Discrete TA (DTA) [74], Continuous TA (CTA) [74], and Open-World TA (OWTA) [53].
@@ II-C Scenarios in CL
CL Difference between $\mathcal{D}_{t-1}$ and $\mathcal{D}_{t}$ Task ID Online
Scenario
p$(\mathcal{X}_{t-1})\neq$ p$(\mathcal{X}_{t})$
p$(\mathcal{Y}_{t-1})\neq$ p$(\mathcal{Y}_{t})$
@@ II-C1 Task-Incremental Learning (Task-IL)
II-C2 Domain-Incremental Learning (Domain-IL)
II-C3 Class-Incremental Learning (Class-IL)
II-C4 Discrete Task Agnostic (DTA)
@@
II-C5 Open-World Task Agnostic (OWTA)
II-C6 Continuous Task Agnostic (CTA)
@@
II-D Online Continual Learning (OCL)
@@ II-E CL Methods
II-E1 Representation-based Approach
@@
Figure 7: An illustration of the taxonomy of (a) representation-based approach and (b) its respective methods; adapted from studies in [16].
  1. 1.
@@
II-E2 Regularization-based Approach
@@
Figure 8: An illustration of the taxonomy of (a) regularization-based approach and (b) its respective methods; adapted from studies in [16].
  1. 1.
@@
II-E3 Rehearsal/Replay-based Approach
Figure 9: An illustration of the taxonomy of (a) replay-based approach and (b) its respective methods; adapted from studies in [16].
@@ II-E3 Rehearsal/Replay-based Approach
  1. 1.
@@
II-E4 Optimization-based Approach
@@
Figure 10: An illustration of the taxonomy of (a) optimization-based approach and (b) its respective methods; adapted from studies in [16].
  1. 1.
@@
II-E5 Architecture-based Approach
@@
Figure 11: An illustration of the taxonomy of (a) architecture-based approach and (b) its respective methods; adapted from studies in [16].
  1. 1.
@@
    Figure 12: An illustration of the model decomposition method exhibiting two types; corresponding to parameters (i.e., low-rank factorization) and representations (i.e., masks of intermediate features); adapted from studies in [16].
All these existing methods for fulfilling the desiderata of CL are also summarized and presented in Table III.
TABLE III: Summary of CL methods based on Representation-, Regularization-, Rehearsal/Replay- and Architecture-based approaches that potentially fulfill “+” or have fulfilled “✓” the desiderata of CL.
Desiderata of CL
Approach Methods Work,(Year)
@@
II-F CL Settings
II-F1 Supervised Learning
@@ II-G Evaluation Metrics
III Neuromorphic Computing
III-A Overview of SNNs
@@
III-B Neuromorphic Continual Learning (NCL)
NCL is an advancement of bio-plausible SNNs for facilitating continual adaptation in dynamic environments. Here, we discuss the state-of-the-art NCL methods, including their key ideas and limitations, whose details are summarized in Table IV.
III-B1 Enhancements on Unsupervised STDP Learning
While conventional STDP enables efficient online unsupervised learning, employing STDP alone may still suffer from CF [41, 42]. To address these challenges, proposed techniques have been proposed in the literature, as discussed in the following.
@@ III-B3 Active Dendrites
III-B4 Bayesian Continual Learning
@@
III-B5 Architecture-based Approach
@@
III-B6 Rehearsal/Replay-based Approach
@@
III-B7 Regularization-based Approach
@@ III-B8 Hebbian Learning
Figure 24: An illustration of the HLOP; adapted from studies in [320]. (a) Overview of the method. (b) Skew-symmetric connection weights in lateral circuits. (c) Construction of new subspaces for new tasks with Hebbian learning in lateral circuits. (d) Recurrent lateral connections based orthogonal projection.
III-C Efficiency Enhancement Methods for NCL
@@ III-C2 Weight Quantization
TABLE IV: Summary of the state-of-the-art works for energy-efficient CL employing bio-plausible SNNs (i.e., NCL).
@@
III-D Summary of NCL
The state-of-the-art methods have performed initial studies for enabling NCL (Table IV). We observe that methods with bio-plausible learning rules (e.g., STDP) facilitate unsupervised learning, which is suitable for OCL scenario. These works mainly employ fully-connected SNN architectures (see Table V), consisting of input layer and a pair of excitatory-inhibitory layer. As the input data grows in complexity, fully-connected SNNs struggle to capture the important hierarchical features without increasing the network size. Larger architectures often lead to better accuracy but require larger storage sizes, which may exceed the resource constraints. This limits the deployable networks on practical hardware platforms, and make their performances limited to simple datasets (e.g., MNIST) with simple CL scenarios (e.g., Split-MNIST). Meanwhile, methods that employ supervised-based learning rules (e.g., surrogate gradient-based backpropagation) can achieve acceptable performance on complex datasets (e.g., ImageNet). However, they are not suitable for OCL scenario, as they require costly labeled data and power-hungry training. Furthermore, these works mainly focus on energy-efficiency, and have not optimized resource utilization. Therefore, alternative methods are required especially for enabling energy-efficient SNNs with OCL capabilities, which are beneficial for many real-world application use-cases.
TABLE V: Comparative analysis of the state-of-the-art in terms of network architecture, model size, performance, memory footprint, and power/energy savings. Note, FC: fully-connected layer, and CONV: convolutional layer.
@@ III-D Summary of NCL
IV Real-World Application Use-Cases
@@
IV-A Adaptive Robots
IV-A1 Object Recognition for Humanoid Robots
@@
IV-A2 Inverse Kinematics and PID Control
@@
IV-B Autonomous Vehicles and Mobile Agents
@@ IV-B1 Cars Detection
IV-B2 Road Lane Detection
Other SNN works for low-power autonomous vehicles include the road line detection. For instance, LaneSNN [348] addressed the imbalance between lane and background classes by employing a novel loss function combining Weighted Binary Cross Entropy (WCE) and Mean Squared Error (MSE); see Fig. 29. Evaluations demonstrated comparable performance to state-of-the-art methods like LaneNet and RefineNet but with significantly lower power consumption. However, it still considered supervised settings which require costly labeled data, power-hungry training, and long training time to update its knowledge, hence it is not suitable for OCL scenario.
@@ IV-B2 Road Lane Detection
IV-B3 Simultaneous Localization and Mapping (SLAM)
@@ IV-B4 People Detection
A summary of real-world application use-cases that will benefit from SNNs with OCL, is presented in Table VI.
TABLE VI: Summary of real-world application use-cases that will benefit from SNNs with OCL capabilities, and their current state-of-the-art works.
Applications Work CL Dataset CL Learning Neuron Hardware
(Year) Setting Approach Rule Model
Object [341] Supervised Custom Representation- Modified LIF Loihi
@@
Autonomous
Vehicles/
Agents SLAM [349] Unsupervised Real-time data Representation- STDP LIF n/a
@@
V Open Research Challenges
@@
Adapting the Existing CL Methods in SNN Domain: Investigating how existing CL methods in Section II-E can be adapted for SNNs is potential for combining their strengths. For instance, adapting the regularization-based methods and how it impacts the performance of SNNs trained with local learning rules, particularly in dynamic environments could be explored. The optimal memory size for replay-based method in SNNs that balances desiderata effectively, particularly in resource-constrained environments can also be explored.